OpenStudy (anonymous):

Show that \[\int_0^\infty \cos(x^2)\,dx=\int_0^\infty \sin(x^2)\,dx=\sqrt{\frac{\pi}{8}}\]

OpenStudy (anonymous):

[ A relatively simple problem compared to the recent ones :) ]

OpenStudy (michele_laino):

I can write the result as below: \[\Im \left( \int\limits_{0}^{+\infty} e ^{ix ^{2}} dx\right)\] then I use this substitution: \[ix ^{2}=y\] and observing that \[\sqrt{i}=\frac{ 1 }{ \sqrt{2} }(1\pm i)\] I got this:\[-\frac{ 1 }{ 2 \sqrt{i} }\Gamma \left( \frac{ 1 }{ 2 } \right)\] so the final result is: \[\pm \frac{ \sqrt{\pi} }{ 2\sqrt{2} }\]

OpenStudy (michele_laino):

where:\[\Im \] stands for imaginary part of a complex number

OpenStudy (michele_laino):

oops...I used: \[ix ^{2}=-y\]

OpenStudy (anonymous):

Well done, I had a similar method, but I keep tripping up with the complex computation: \[\begin{align*} \int_{-\infty}^\infty e^{-x^2}\,dx&=\int_{-\infty}^\infty e^{(ix)^2}\,dx\\\\ \sqrt\pi&=\int_{-\infty}^\infty\bigg(\cos (ix)^2+i\sin(ix)^2\bigg)\,dx\\\\ &=\int_{-\infty}^\infty\cos(x^2)\,dx-\underbrace{\mathrm{Im}\left[\int_{-\infty}^\infty e^{-ix^2}\,dx\right]}\end{align*}\] The bracketed part is the part that's particularly bothersome. At any rate, once I figure this out I'd divide by 2 and the result follows.

OpenStudy (anonymous):

I get that the bracketed part amounts to \(-\dfrac{\sqrt{\pi}}{\sqrt2}\), so I suppose there's a mistake in my reasoning somewhere along the line...

OpenStudy (anonymous):

It's probably due to \[\large\int_\mathbb{R}e^{-ax^2}\,dx=\sqrt{\frac{\pi}{a}}\quad\text{for Re}(a)>0\] and not for complex \(a\).

OpenStudy (michele_laino):

you are right! @SithsAndGiggles

OpenStudy (anonymous):

*not for \(\text{Re}(a)\le0\), I mean

OpenStudy (michele_laino):

I think so!

OpenStudy (michele_laino):

very nice question :) @SithsAndGiggles

OpenStudy (anonymous):

Thanks! I originally considered approaching the integral via differentiation under the integral sign but I couldn't set up the proper parameterization.

OpenStudy (michele_laino):

I understand

OpenStudy (michele_laino):

I see

OpenStudy (michele_laino):

thanks! for your website @SithsAndGiggles